OpenStudy (anonymous):

How do I find the domain of f(x) = (-x^2+16/x^2-2x-3)

OpenStudy (anonymous):

whenever you have a fraction function, the domain will be: (bottom part) > 0

OpenStudy (anonymous):

false; it is all x values that will make the bottom not equal to 0.

OpenStudy (anonymous):

so, \[x^{2} - 2x - 3 \neq 0 \] solve for x and it'll give yu yur domain

OpenStudy (anonymous):

set the denominator = 0 solve. x^2-2x-3=0 (x+1)(x-3)=0 x=-1 or x = 3 domain is all real numbers except -1 and 3